# TME-Box: Scalable In-Process Isolation through Intel TME-MK Memory Encryption

### v1

@@ I Introduction
Cloud computing must be highly optimized for performance and efficiency, processing requests with the lowest possible latencies. Various techniques and optimizations are implemented at the software and infrastructural level of the cloud architecture. Common optimizations are the exclusion of protection mechanisms, like process isolation, in favor of low start-up times and fast execution for serverless applications in a multi-tenant environment, e.g., function-as-a-service (FaaS) applications [22, 20, 14, 33, 73]. However, omitting heavy-weight protection mechanisms such as process isolation introduces substantial security risks. A single memory safety vulnerability [54, 27] can result in a full compromise of the entire cloud computing system. Consequently, attackers can compromise other tenant’s highly sensitive data.
@@
The efficiency of isolation can be significantly enhanced with hardware support. Hardware-based isolation mechanisms, such as protection keys for user space (PKU) [61], enable the restriction of access to memory at a page granularity. Specifically, memory protection keys (MPK) use a 4-bit protection key located in the page table entries (PTE) to assign and enforce access policies for the respective page. These policies are software-controlled via a processor register, associating two bits for every protection key to dynamically control read and write permissions. Thereby, MPK allows the partitioning of the process memory into 16 distinct regions, often called domains, and selectively controls their access. Various PKU-based protection schemes [81, 68, 62, 7] are proposed to establish lightweight and transparent in-process isolation that is efficiently enforced in hardware. However, MPK is limited by its 4-bit key size, only offering 16 distinct protection domains. This is restrictive for software systems that may require thousands of concurrent domains, e.g., Cloudflare Workers [14, 17], within a single process. Also, MPK is bound to the page granularity and cannot provide sub-page granular isolation.
@@
In this paper, we present TME-Box, a novel sandboxing technique for scalable in-process isolation that repurposes Intel’s newly introduced total memory encryption multi-key (TME-MK) [34] technology. TME-MK’s runtime memory encryption is essentially used by the Intel trusted domain extensions (TDX) [35], their confidential computing architecture, to ensure the confidentiality and integrity of memory resources. This encryption and other architectural elements (e.g., secure extended page tables (EPT) [39]) allow TDX to provide strong isolation of virtual machines (VMs), which run isolated workloads on the same physical machine.
@@
TME-Box extends the application of TME-MK beyond hardware-isolated VMs by providing fine-grained and scalable in-process isolation on commodity x86 CPUs. TME-Box deploys compiler instrumentation to isolate the code and data of mutually untrusted sandboxes. Specifically, we enforce the usage of sandbox-specific TME-MK encryption keys by controlling the base address and index of memory operations. Thereby, memory resources of the in-process sandbox are cryptographically isolated, leading to the detection of unauthorized memory access through hardware-backed integrity protection. Repurposing the TME-MK memory encryption for sandboxing offers several advantages over established in-process isolation mechanisms and addresses the hard requirements of modern computing systems, such as cloud computing.
@@
Second, TME-Box supports the flexible relocation of data in memory, which is particularly important for data-centric computation. This enables flexible memory management for the allocator, providing continuous memory utilization across different sandboxes and efficient memory migration.
@@
TME-Box. We are the first to present a novel in-process isolation technique by repurposing the Intel TME-MK encryption on commodity x86 CPUs.
@@ II-A Address Translation
In modern operating systems, virtual memory organizes memory into contiguous blocks known as memory pages, which are typically in size on the x86-64 architecture. Pages are managed in a hierarchical structure called page tables, enabling each user space application to have a distinct virtual representation of the computer’s physical memory. Address translation in paging systems is performed by the memory management unit (MMU), which is responsible for resolving the virtual-to-physical mappings. When accessing memory, the CPU translates a virtual address to a physical address via a page table walk. The MMU uses the corresponding page table entries (PTE) that contain the physical page number (PPN) (and associated access permissions) to resolve the virtual address. Frequently used virtual-to-physical mappings are cached in the translation look-aside buffer (TLB) to enhance system performance.
@@ II-B Software-based Fault Isolation
Memory safety vulnerabilities [77] frequently occur in complex software written in memory-unsafe programming languages (e.g., C, C++) [54, 27]. This is critical for software security, as an adversary can exploit memory safety vulnerabilities to compromise the target system [2, 63]. For instance, out-of-bounds (OOB) errors, such as buffer over-reads and over-writes, enable an adversary to illegally access resources in memory. The exploitation of memory safety errors allows the attacker to leak or corrupt sensitive data. To mitigate this security threat, software sandboxing is used. Sandboxes aim to reduce the attack surface of memory errors by isolating individual parts of the software.
Software-based fault isolation (SFI) [82] is a defense technique that implements in-process isolation (also referred to as intra-process or sub-process isolation), separating memory resources for individual software components within a single process. Each SFI sandbox consists of an isolated data region containing runtime data (e.g., stack and heap memory) and an isolated code region where the code of the sandbox resides [78]. Consequently, the SFI sandbox needs to restrict data access and control-flow transfers. Thus, all memory operations (i.e., read and write accesses) performed by the sandbox’s code are only allowed to access data within this isolated data region. In addition, control-flow transfers must remain within the sandbox’s code region (or target call sites that correspond to trusted runtime calls) [78].
@@ II-C Intel Defensive Execution Technologies
Intel Control-Flow Enforcement Technologies
Control-flow enforcement technologies (CET) [76] is a set of processor extensions designed to implement control-flow integrity (CFI) [1, 10, 11] measures into the CPU hardware. CFI can be loosely categorized into two types: forward-edge and backward-edge CFI, which aim to protect function pointers and return addresses, respectively.
In the context of forward-edge CFI, Intel CET integrates indirect branch tracking (IBT), utilizing so-called landing pads identified by the endbr64 instructions inserted into function entries by the compiler. Valid indirect jump targets are limited to function entries, thus reducing the attack surface of control-flow hijacking attacks such as jump-oriented programming (JOP) [8]. In terms of backward-edge CFI, Intel CET integrates a hardware shadow stack feature, which applies protection for return addresses. The shadow stack feature achieves this by copying the return address onto the shadow stack, which is inaccessible to an attacker when entering a function. On function exit, the return address restored from the regular stack is compared against the one on the shadow stack. Mismatches are detected in hardware, and the potential corruption of return addresses is mitigated. Thereby, CET thwarts code-reuse attacks such as return-into-libc and return-oriented programming (ROP) [74].
@@
Figure 2 illustrates the encryption procedure of TME-MK with integrity support. The TME-MK encryption engine uses AES [19, 18] in XTS mode [50, 64], supporting both 128-bit and 256-bit encryption keys. In addition, with the introduction of the Intel trust domain extensions (TDX) [35], authenticated encryption is integrated. This extends TME-MK with support for data integrity using a message authentication code (MAC). The MAC is computed using the SHA-3 [30] secure hash algorithm. Memory operations are performed using the associated keyID that selects the used encryption key. When writing to memory, TME-MK encrypts the data and computes the MAC to provide confidentiality and integrity of DRAM data (cf. Figure 2(a) - Memory Encryption). When reading encrypted data from DRAM, it decrypts and authenticates the data by recomputing and comparing the associated MAC. An integrity violation is detected within the cryptographic bounds of the MAC (cf. Figure 2(b) - Memory Decryption). In case of an integrity violation, either a fixed pattern is returned to prevent ciphertext analysis [35], or a hardware exception is raised [66].
@@ IV-A Secure System Architecture
In the following, we outline the main design properties of TME-Box, addressing the hard requirements of modern computing systems, such as cloud computing. We enable scalable isolation and flexible data relocation without introducing new software dependencies, e.g., hard constraints for the memory allocator. We provide cryptographic in-process isolation, thereby supporting a large number of sandboxes (e.g., for isolating cloud workers) and detecting unauthorized access through integrity exceptions. Thereby, TME-Box enables hardware-supported in-process isolation on commodity x86 machines for modern cloud settings.
@@ IV-B Scalable Memory Isolation and Flexible Data Relocation
Our scheme enables the flexible relocation of data located in memory, which is important for software systems that focus on data-centric computation. This allows the memory allocator to efficiently migrate memory, thereby reducing the overall memory fragmentation and consumption. Figure 5 provides an overview of TME-Box data relocation. The figure shows the relocation of sparsely allocated cache lines from different sandboxes on memory pages A and B to a single page, page C. This results in more efficient management of fragmented memory resources.
@@ IV-C Isolation of In-Process Sandboxes
TME-Box repurposes Intel TME-MK to achieve scalable memory isolation. Any violation of TME-Box’s memory access policies, i.e., one sandbox trying to access another sandbox’s memory, is detected by the TME-MK encryption engine. TME-Box relies on TME-MK’s integrity checks to detect sandbox access violations. In particular, TME-MK associates a cryptographic MAC that is 28-bit in size with each cache line [35]. Memory is initialized by writing the entire cache line, which sets up the MAC with the used keyID, thus providing data integrity. For subsequent memory reads, TME-MK authenticates the MAC using the stored DRAM data and keyID of the access.
TME-Box leverages this authentication procedure to detect unauthorized memory accesses of sandboxes. Memory requests with an incorrect keyID result in a MAC mismatch, triggering an exception. Similarly, (partial) writes to a cache line with an incorrect keyID corrupts the MAC associated with the cache line. In this case, the access violation is detected when the sandbox owning the memory location performs a read access. This allows TME-Box to detect access violations within the security bounds provided by the cryptographic MAC.
@@
For forward-edge transfers, TME-Box instruments indirect function calls and jumps (i.e., function pointer dereferences) to stay within the sandbox’s isolated code region. Address masking is applied to these function pointers, enforcing that code execution is confined within the sandbox. Similarly, backward-edge transfers, such as function returns, are transformed and instrumented to ensure that return addresses target code within the sandbox. Additionally, direct function calls are analyzed during compile-time to ensure that they target valid call sites, i.e., authorized call targets within the sandbox or trusted runtime calls.
@@ V-A Compiler Extension
We base our prototype implementation on the LLVM compiler infrastructure [47] (version 14.0.0). Our compiler extension consists of a set of individual compiler passes required for TME-Box’s isolation of code and data. Moreover, the TME-Box framework is fully parameterizable regarding the utilized virtual address space (e.g., 57-bit virtual addressing). We also detail architecture-specific optimizations in our prototype implementation, resulting in more efficient code instrumentation. Specifically, TME-Box integrates the following three compiler modifications:
@@
First, we need to reserve a dedicated CPU register to securely store the base address of the sandbox and enable fast access. TME-Box implements two options for this: We either apply an x86-specific optimization by using a segment register222The x86 architecture supports segmentation, allowing memory instructions to use a segment-based addressing mode. In particular, the fs and gs segment registers are still functional in 64-bit mode. While the fs segment is commonly used to address thread local storage (TLS), the gs segment has no common use and, thus, can be utilized by applications [42]. [42] with segment-relative memory accesses or reserve a general-purpose register. Our compiler framework supports both options, allowing for maximal compatibility and performance comparison, which is reflected in our evaluation. Depending on the chosen TME-Box mode, we either reserve the gs segment register or a single general-purpose register, in our case r15, to store the sandbox’s base address (i.e., the upper part of the virtual address) that corresponds to the sandbox’s assigned keyID in the PTEs, which maps to the designated encryption key of the sandbox. Note that reserving a general-purpose CPU register has performance implications due to increased register pressure. Future TME-Box implementations can use the Intel advanced performance extensions (APX)333Intel Advanced Performance Extensions (APX) [37] enhance x86 processors by increasing the number of general-purpose registers from 16 to 32. [37] that enhance x86 processors with an increased register set. This optimization would then also lessen the overhead of reserving a general-purpose register.
@@
Lastly, we implement a compiler pass that isolates the code of the sandbox by instrumenting control-flow transfers. Precisely, forward edge and backward edge transitions are instrumented via address masking, restricting transfers to remain in the sandbox’s code region, comparable to recent SFI-based sandboxing [86, 69]. For forward edge transfers (i.e., indirect function calls and jumps), the compiler instruments the corresponding function pointers before executing the dedicated instruction. This restricts addressable call sites to the virtual address space of the respective sandbox. Similarly, backward edge transfers (i.e., function returns) are instrumented. Here, our compiler transforms returns into a code sequence that receives the return address from the stack, instruments it, and performs an indirect jump. As an optimization, a future TME-Box implementation can use Intel CET555Intel Control-flow Enforcement Technologies (CET) [76] integrates a shadow stack feature that applies return address protection in hardware. [76] for return address protection, increasing security and enhancing system performance. Additionally, direct function calls are checked on compile-time for valid call sites, i.e., the sandbox is allowed to call this function directly.

### v2

@@ I Introduction
Cloud computing must be highly optimized for performance and efficiency, processing requests with the lowest possible latencies. Various techniques and optimizations are implemented at the software and infrastructural level of the cloud architecture. Common optimizations are the exclusion of protection mechanisms, like process isolation, in favor of low start-up times and fast execution for serverless applications in a multi-tenant environment, e.g., function-as-a-service (FaaS) applications [22, 20, 16]. However, omitting heavy-weight protection mechanisms such as process isolation introduces substantial security risks. A single memory safety vulnerability [54, 82] can result in a full compromise of the entire cloud computing system. Consequently, attackers can compromise other tenant’s highly sensitive data.
@@
The efficiency of isolation can be significantly enhanced with hardware support. Hardware-based isolation mechanisms, such as protection keys for user space (PKU) [61], enable the restriction of access to memory at a page granularity. Specifically, memory protection keys (MPK) use a 4-bit protection key located in the page table entries (PTE) to assign and enforce access policies for the respective page. These policies are software-controlled via a processor register, associating two bits for every protection key to dynamically control read and write permissions. Thereby, MPK allows the partitioning of the process memory into 16 distinct regions, often called domains, and selectively controls their access. Various PKU-based protection schemes [84, 72, 71, 31, 8] are proposed to establish lightweight and transparent in-process isolation that is efficiently enforced in hardware. However, MPK is limited by its 4-bit key size, only offering 16 distinct protection domains. This is restrictive for software systems that may require thousands of concurrent domains, e.g., Cloudflare Workers [16, 17], within a single process. Also, MPK is bound to the page granularity and cannot provide sub-page granular isolation.
@@
In this paper, we present TME-Box, a novel sandboxing technique for scalable in-process isolation that repurposes Intel’s newly introduced total memory encryption multi-key (TME-MK) [32] technology. TME-MK’s runtime memory encryption is essentially used by the Intel trusted domain extensions (TDX) [15, 33], their confidential computing architecture, to ensure the confidentiality and integrity of memory resources. This encryption and other architectural elements (e.g., secure extended page tables (EPT) [36]) allow TDX to provide strong isolation of virtual machines (VMs), which run isolated workloads on the same physical machine.
@@
TME-Box extends the application of the TME-MK memory encryption beyond hardware-isolated VMs by providing fine-grained and scalable in-process isolation on commodity x86 CPUs. TME-Box deploys compiler instrumentation to isolate the code and data of mutually untrusted sandboxes. Specifically, we enforce the usage of sandbox-specific TME-MK encryption keys by controlling the base address and index of memory operations. Thereby, memory resources of the in-process sandbox are cryptographically isolated, leading to the detection of unauthorized memory access through hardware-backed integrity protection. Repurposing the TME-MK memory encryption for software sandboxing offers several advantages over established in-process isolation mechanisms and addresses the hard requirements of modern computing systems, such as cloud computing.
@@
Second, TME-Box supports the flexible relocation of data in memory, which is particularly important for data-centric computation. Data relocation is essential in the cloud, e.g., when allocator caches with smaller heap slots are used. This enables flexible memory management for the allocator, providing continuous memory utilization across different sandboxes and efficient memory migration.
@@
TME-Box. We are the first to present a novel in-process isolation technique by repurposing the Intel TME-MK memory encryption on commodity x86 CPUs.
@@ II-A Address Translation
In modern operating systems, virtual memory organizes memory into contiguous blocks known as memory pages, which are typically in size on the x86-64 architecture. Pages are managed in a hierarchical structure called page tables, enabling each user space application to have a distinct virtual representation of the computer’s physical memory.
Address translation in paging systems is performed by the memory management unit (MMU), which is responsible for resolving the virtual-to-physical mappings. When accessing memory, the CPU translates a virtual address to a physical address via a page table walk. The MMU uses the corresponding page table entries (PTE) that contain the physical page number (PPN) (and associated access permissions) to resolve the virtual address. Frequently used virtual-to-physical mappings are cached in the translation look-aside buffer (TLB) to enhance system performance.
@@ II-B Software-based Fault Isolation
Memory safety vulnerabilities [80] frequently occur in complex software written in memory-unsafe programming languages (e.g., C, C++) [54, 82]. This is critical for software security, as an adversary can exploit memory safety vulnerabilities to compromise the target system [2, 65]. For instance, out-of-bounds (OOB) errors, such as buffer over-reads and over-writes, enable an adversary to illegally access resources in memory. The exploitation of memory safety errors allows the attacker to leak or corrupt sensitive data.
To mitigate this security threat, software sandboxing is used. Sandboxes aim to reduce the attack surface of memory errors by isolating individual parts of the software. Software-based fault isolation (SFI) [85] is a defense technique that implements in-process isolation (also referred to as intra-process or sub-process isolation), separating memory resources for individual software components within a single process. Each SFI sandbox consists of an isolated data region containing runtime data (e.g., stack and heap memory) and an isolated code region where the code of the sandbox resides [81]. Consequently, the SFI sandbox needs to restrict data access and control-flow transfers. Thus, all memory operations (i.e., read and write accesses) performed by the sandbox’s code are only allowed to access data within this isolated data region. In addition, control-flow transfers must remain within the sandbox’s code region (or target call sites that correspond to trusted runtime calls) [81].
@@ II-C Intel Defensive Execution Technologies
Intel Control-Flow Enforcement Technology
Control-flow enforcement technology (CET) [79] is a set of processor extensions designed to implement control-flow integrity (CFI) [1, 12, 13] measures into the CPU hardware. CFI can be loosely categorized into two types: forward-edge and backward-edge CFI, which aim to protect function pointers and return addresses, respectively.
In the context of forward-edge CFI, Intel CET integrates indirect branch tracking (IBT), utilizing so-called landing pads identified by the endbr64 instructions inserted into function entries by the compiler. Valid indirect jump targets are limited to function entries, thus reducing the attack surface of control-flow hijacking attacks such as jump-oriented programming (JOP) [9]. In terms of backward-edge CFI, Intel CET integrates a hardware shadow stack feature, which applies protection for return addresses. The shadow stack feature achieves this by copying the return address onto the shadow stack, which is inaccessible to an attacker when entering a function. On function exit, the return address restored from the regular stack is compared against the one on the shadow stack. Mismatches are detected in hardware, and the potential corruption of return addresses is mitigated. Thereby, CET thwarts code-reuse attacks such as return-into-libc [77] and return-oriented programming (ROP) [11].
@@
Figure 2 illustrates the encryption procedure of TME-MK with integrity support. The TME-MK encryption engine uses AES [19, 18] in XTS mode [50, 67], supporting both 128-bit and 256-bit encryption keys. In addition, with the introduction of the Intel trust domain extensions (TDX) [15, 33], authenticated encryption is integrated. This extends TME-MK with support for data integrity using a message authentication code (MAC). The MAC is computed using the SHA-3 [7] secure hash algorithm. Memory operations are performed using the associated keyID that selects the used encryption key. When writing to memory, TME-MK encrypts the data and computes the MAC to provide confidentiality and integrity of DRAM data (cf. Figure 2(a) - Memory Encryption). When reading encrypted data from DRAM, it decrypts and authenticates the data by recomputing and comparing the associated MAC. An integrity violation is detected within the cryptographic bounds of the MAC (cf. Figure 2(b) - Memory Decryption). In case of an integrity violation, either a fixed pattern is returned to prevent ciphertext analysis [33], or a hardware exception is raised [70].
@@ IV-A Secure System Architecture
In the following, we outline the main design properties of TME-Box, addressing the hard requirements of modern computing systems, such as cloud computing. We enable scalable isolation and flexible data relocation without introducing new software dependencies, e.g., hard constraints for the memory allocator. We provide cryptographic in-process isolation, thereby supporting a large number of sandboxes (e.g., for isolating cloud workers) and detecting unauthorized access through integrity exceptions. In this way, TME-Box enables hardware-supported in-process isolation on commodity x86 machines for modern cloud settings.
@@ IV-B Scalable Memory Isolation and Flexible Data Relocation
Our scheme enables the flexible relocation of data located in memory, which is important for software systems that focus on data-centric computation. This allows the memory allocator to efficiently migrate memory, thereby reducing the overall memory fragmentation and consumption. Figure 5 provides an overview of TME-Box data relocation. The figure shows the relocation of sparsely allocated cache lines from different sandboxes on memory pages A and B to a single page, page C. This allows the usage of allocator caches and results in a more efficient management of fragmented memory resources.
@@ IV-C Isolation of In-Process Sandboxes
TME-Box repurposes Intel TME-MK to achieve scalable memory isolation. Any violation of TME-Box’s memory access policies, i.e., one sandbox trying to access another sandbox’s memory, is detected by the TME-MK encryption engine. TME-Box relies on TME-MK’s integrity checks to detect sandbox access violations. In particular, TME-MK associates a cryptographic MAC that is 28-bit in size with each cache line [33].
Memory is initialized by writing the entire cache line, which sets up the MAC with the used keyID, thus providing data integrity. For subsequent memory reads, TME-MK authenticates the MAC using the stored DRAM data and keyID of the access. TME-Box leverages this authentication procedure to detect unauthorized memory accesses of sandboxes.
Memory requests with an incorrect keyID result in a MAC mismatch, triggering an exception. Similarly, (partial) writes to a cache line with an incorrect keyID corrupts the MAC associated with the cache line. In this case, the access violation is detected when the sandbox owning the memory location performs a read access. This allows TME-Box to detect access violations within the security bounds provided by the cryptographic MAC.
@@
For forward-edge transfers, TME-Box instruments indirect function calls and jumps (i.e., function pointer dereferences) to stay within the sandbox’s isolated code region. Address masking is applied to these function pointers, enforcing that code execution is confined within the sandbox.
Similarly, backward-edge transfers, such as function returns, are transformed and instrumented to ensure that return addresses target code within the sandbox. Additionally, direct function calls are analyzed during compile-time to ensure that they target valid call sites, i.e., authorized call targets within the sandbox or trusted runtime calls.
@@ V-A Compiler Extension
We base our prototype implementation on the LLVM compiler infrastructure [45] (version 14.0.0). Our compiler extension consists of a set of individual compiler passes required for TME-Box’s isolation of code and data. Moreover, the TME-Box framework is fully parameterizable regarding the utilized virtual address space (e.g., 57-bit virtual addressing).
We also detail architecture-specific optimizations in our prototype implementation, resulting in more efficient code instrumentation. Specifically, TME-Box integrates the following three compiler modifications:
@@
First, we need to reserve a dedicated CPU register to securely store the base address of the sandbox and enable fast access. TME-Box implements two options for this: We either apply an x86-specific optimization by using a segment register222The x86 architecture supports segmentation, allowing memory instructions to use a segment-based addressing mode. In particular, the fs and gs segment registers are still functional in 64-bit mode. While the fs segment is commonly used to address thread local storage (TLS), the gs segment has no common use and, thus, can be utilized by applications [39]. [39] with segment-relative memory accesses or reserve a general-purpose register. Our compiler framework supports both options, allowing for maximal compatibility and performance comparison, which is reflected in our evaluation.
Depending on the chosen TME-Box mode, we either reserve the gs segment register or a single general-purpose register, in our case r15, to store the sandbox’s base address (i.e., the upper part of the virtual address) that corresponds to the sandbox’s assigned keyID in the PTEs, which maps to the designated encryption key of the sandbox. Note that reserving a general-purpose CPU register has performance implications due to increased register pressure. Future TME-Box implementations can use the Intel advanced performance extensions (APX)333Intel Advanced Performance Extensions (APX) [37] enhance x86 processors by increasing the number of general-purpose registers from 16 to 32. [37] that enhance x86 processors with an increased register set. This optimization would then also lessen the overhead of reserving a general-purpose register.
@@
Lastly, we implement a compiler pass that isolates the code of the sandbox by instrumenting control-flow transfers. Precisely, forward-edge and backward-edge transitions are instrumented via address masking, restricting transfers to remain in the sandbox’s code region, comparable to recent SFI-based sandboxing [89, 73].
For forward-edge transfers (i.e., indirect function calls and jumps), the compiler instruments the corresponding function pointers before executing the dedicated instruction. This restricts addressable call sites to the virtual address space of the respective sandbox. Similarly, backward-edge transfers (i.e., function returns) are instrumented. Here, our compiler transforms returns into a code sequence that receives the return address from the stack, instruments it, and performs an indirect jump. As an optimization, a future TME-Box implementation can use Intel CET555Intel Control-flow Enforcement Technology (CET) [79] integrates a shadow stack feature that applies return address protection in hardware. [79] for return address protection, increasing security and enhancing system performance. Additionally, direct function calls are checked on compile-time for valid call sites, i.e., the sandbox is allowed to call this function directly.